# A unified framework for equation discovery and dynamic prediction of hysteretic systems

### v1

@@ Abstract
Hysteresis is a nonlinear phenomenon with memory effects, where a system’s output depends on both its current state and past states. It is prevalent in various physical and mechanical systems, such as yielding structures under seismic excitation, ferromagnetic materials, and piezoelectric actuators. Analytical models like the Bouc-Wen model are often employed but rely on idealized assumptions and careful parameter calibration, limiting their applicability to diverse or mechanism-unknown behaviors. Existing equation discovery approaches for hysteresis are often system-specific or rely on predefined model libraries, which limit their flexibility and ability to capture the hidden mechanisms. To address these, this research develops a unified framework that integrates learning of internal variables (commonly used in modeling hysteresis) and symbolic regression to automatically extract internal hysteretic variable, and discover explicit governing equations directly from data without predefined libraries as required by methods such as sparse identification of nonlinear dynamics (SINDy). Solving the discovered equations naturally enables prediction of the dynamic responses of hysteretic systems. This work provides a systematic view and approach for both equation discovery and characterization of hysteretic dynamics, defining a unified framework for these types of problems.
@@
Hysteretic dynamic systems; equation discovery; symbolic regression; internal variable learning
@@ 1 Introduction
Hysteresis is a common nonlinear phenomenon observed in a broad spectrum of engineering and physical systems, including ferromagnetic materials [jiles1986theory], piezoelectric actuators [al2008modeling], structural materials [bertotti1998hysteresis], damping devices [ko1999design, dominguez2004modelling], and shape memory alloys [ortin2002hysteresis]. Hysteresis, characterized by its path-dependent behavior, refers to systems in which the output depends not only on the current state, but also on the history of past states. This memory effect often appears as input-output loops, and is especially evident in systems that involve energy dissipation, plastic deformation, or phase transitions [wen1976method, visone2008hysteresis]. In nonlinear dynamic systems, hysteresis presents significant challenges for modeling, control, and prediction, owing to its multivalued behavior, rate dependence, and sensitivity to input history. Accurately capturing hysteresis behavior is therefore crucial in applications such as structural health monitoring [chatzi2010experimental], smart material design [chandra2023discovery], and precision actuation systems [liu2022experimental, vlachas2021two].
In response to these needs, numerous mathematical models have been proposed over the past decades to characterize hysteretic behaviors. Among these, the Bouc-Wen model [wen1976method] is one of the most widely adopted, owing to its capability to replicate a wide variety of hysteresis loops via differential equations that describe the evolution of an internal variable. Its parameters provide physical interpretability and offer flexibility in capturing both softening and hardening behaviors. In addition to the Bouc-Wen model, other approaches such as the Preisach [mayergoyz2003mathematical], Duhem [oh2005semilinear], and Prandtl-Ishlinskii [visintin2013differential] models offer complementary frameworks, and have been effectively applied in domains where rate-independent behavior or saturation effects are predominant. However, these models typically require prior knowledge of the system structure or involve challenging calibration procedures, limiting their generalizability to complex or mechanism-unknown hysteretic dynamics.
@@
In recent years, data-driven modeling has opened new venues for discovering governing laws directly from measured data [rudy2017data, CHEN2025118330, haywood2024discussing, YI2025112149, ijcai2025p879, yan2025deep]. The sparse identification of nonlinear dynamics (SINDy) [brunton2016discovering] has shown strong potential in extracting compact and interpretable dynamic equations, by applying sparse regression with a predefined library of candidate functions. Extensions such as SINDy with control [brunton2016discovering], physics-informed SINDy [champion2019discovery, chen2021physics], implicit SINDy [kaheman2020sindy], and a sparse structural system identification method applied to hysteretic systems [lai2019sparse] have significantly expanded its applicability to control systems, hidden-variable models, constrained physical systems, and hysteretic dynamics (our focus in this paper). Nevertheless, the effectiveness of SINDy-type methods strongly relies on the expressiveness of the chosen function library, which constrains their ability to capture non-polynomial dynamics, rate-dependent behaviors, and discontinuities. Selecting an appropriate function library is itself a nontrivial task, often requiring prior knowledge or trial-and-error tuning. Furthermore, many dynamical systems involve essential internal states—referred to as internal hysteretic variables in hysteretic systems, or simply as internal variables for brevity in this paper – which are not directly observable but are required to capture implicit evolution and memory effects. Consequently, these methods are not well-suited for such cases.
These limitations become particularly critical in the context of hysteretic systems, where complex memory-dependent behaviors introduce additional challenges. (i) First, internal variables – such as restoring forces or internal displacements – are often not directly measurable [noel2016hysteretic], complicating the construction of state-space representations. (ii) Second, hysteresis is typically characterized by strong memory and rate dependence, which static or memoryless models are unable to capture. In this research, the proposed framework addresses these issues through a proposed solver-based internal variable learning scheme, which enables the learning of hysteretic variables (see Section 3.2 for details). (iii) Third, the presence of non-smooth mathematical operations, such as absolute values or sign functions, precludes the use of Taylor expansions and hinders the application of gradient-based optimization techniques. Our framework further tackles this challenge by employing equation discovery to flexibly utilize data for recovering explicit expressions of complex structures (see Section 3.3 for details). These challenges become even more pronounced in dynamic systems where the governing laws are implicit or nonlinear in unstructured ways. Consequently, there is a growing need for modeling frameworks capable of simultaneously inferring internal variables, capturing memory effects, and producing interpretable governing equations directly from data.
@@
Figure 1: Classification of equation discovery approaches for hysteretic systems according to the assumed knowledge of the primary dynamics motion equation ($f_{\theta}$) and the hysteretic link equation ($g_{\phi}$). The state variables $x,\dot{x},\ddot{x},z\in\mathbb{R}^{1}$ denote the scalar displacement, velocity, acceleration, and internal hysteretic variable, respectively. $\mathbf{y}\in\mathbb{R}^{d}$ denotes the available measurements with selection matrix $\mathbf{S}$ specifying the observed quantities. All these variables are dependent on $t$, which is omitted here for brevity. Each quadrant represents a level of prior structural specification.
@@
A variety of approaches have been proposed for equation discovery in hysteretic systems, which in this paper are classified according to the extent of prior knowledge assumed about the governing equations. In Figure 1, we organize the literature into four categories based on whether the structure of the primary dynamics equation ($f_{\theta}$) and the hysteretic link equation ($g_{\phi}$) are explicitly known or unknown. In the majority of existing studies [lai2019sparse, bolourchi2015studies], structural dynamics is assumed to follow an explicit linear form or other types of known forms, while the hysteretic component is represented by an additional differential equation whose functional forms are to be discovered. Under this setting, the equation discovery task is reduced to symbolically identifying the hysteretic link equation (Hysteresis Discovery quadrant) [wang2025symbolic, liu2025structured] or estimating parameters when structures of both equations are known (Parameter Estimation quadrant) [lin2022identification, worden2012parameter, charalampakis2008identification]. Such assumptions greatly simplify the problem but restrict applicability to systems where the underlying structural model is already well characterized. In contrast, frameworks that assume unknown (implicit) structures in the primary dynamics equation (Dynamics Discovery quadrant) are far less common, particularly when hysteretic effects are present. When both the structural dynamics and the hysteretic link equations are unknown (Full Equation Discovery quadrant), the problem becomes substantially more challenging, requiring the simultaneous inference of two coupled, potentially nonlinear and memory-dependent governing equations. To the best of our knowledge, no prior work has demonstrated a fully symbolic equation discovery framework capable of addressing this most general setting. This highlights the need for a unified approach that can infer both structural and hysteretic dynamics directly from data, without relying on predefined equation forms.
@@
In this research, we propose a unified framework that integrates internal variable learning and symbolic regression (SR) to address the challenges of hysteretic equation discovery. The framework enables direct inference of internal hysteretic variables and recovery of explicit and interpretable governing equations without relying on predefined model libraries. Our contributions are threefold: (i) we provide a classification of equation discovery approaches for hysteretic systems and establish a unified framework that accommodates different levels of prior knowledge; (ii) we introduce a solver-based scheme that enables learning of the internal hysteretic variable and a SR strategy that discovers nonlinear and fractional hysteretic laws directly from data; (iii) we validate the framework on synthetic benchmarks, a complex structure with high-order nonlinear and fractional terms, and shake-table experiments, demonstrating superior accuracy and generalizability compared with existing methods.
@@ 2.1 Hysteretic models
In modeling hysteretic systems, internal variables are commonly introduced to mathematically represent memory effects and capture nonlinear, history-dependent behavior, which allows the system’s response to account for past states.
Among hysteresis models, the Bouc-Wen model [wen1976method] is one of the most widely used, owing to its simplicity, and its capacity to represent a broad range of hysteretic responses. It provides a set of differential equations to describe the evolution of internal variables, and is capable of capturing both rate-dependent and asymmetric behaviors. The dynamics motion equation for a single degree-of-freedom (SDOF) nonlinear mass-spring-damper system with hysteresis can be expressed as follows:
@@
where $k$ is the stiffness, $\alpha$ is the hysteretic coefficient, and $z(t)$ is the internal hysteretic variable. Substituting equation Eq.(2) into Eq.(1), the equation of motion is rewritten as:
@@
The evolution of the internal hysteretic variable $z(t)$ is governed by the Bouc-Wen hysteresis differential equation, which can also be written as an explicit function of $\dot{x}(t)$ and $z(t)$ for compactness and interpretability (for conciseness, we drop $t$ in the equation):
@@
In this formulation, the parameter $A$ serves as a linear amplification factor that scales the influence of the displacement rate on the evolution of the internal variable. The coefficients $\beta$ and $\gamma$ jointly shape the energy dissipation and curvature of the hysteresis loop, with $\beta$ primarily influencing the width and dissipation of the loop, while $\gamma$ controls its nonlinear curvature. In most existing research [qian2024discovering, lai2019sparse, liu2025structured], the exponent $n$ is typically restricted to integer values, such as $n=1$. In this research, we generalize the formulation by allowing $n$ to take more general, including fractional values, thereby enabling richer characterization of various hysteresis.
The variable $z(t)$, representing the internal hysteretic variable, is typically unmeasurable in practical scenarios, and must be inferred from the dynamics of the observed system. This unobservability poses significant challenges for equation discovery, and motivates the use of learning-based approaches to recover the governing evolution law of $z(t)$ from data. Traditional approaches often rely on optimization-based techniques [charalampakis2010identification, noel2017nonlinear], which may suffer from issues such as non-convexity and sensitivity to noise, while other methods are built upon restrictive assumptions or depend on predefined functional libraries, such as SINDy [novelli2025identification, chandra2023discovery, brunton2016discovering]. Moreover, regression-based formulations like SINDy typically do not impose temporal coherence across time series, which limits their ability to capture complex hysteretic behaviors. More recently, physics-guided machine learning methods have been introduced to address these limitations [liu2021physics, liu2022physics], including hybrid models that integrate neural networks with the physical structure of Bouc-Wen dynamics. These approaches aim to preserve physical interpretability while harnessing the representational power of data-driven models. It is worth noting that the Bouc-Wen model here serves merely as a representative example, while the proposed framework is not restricted to this specific formulation, and can be applied to a broader class of hysteretic systems.
@@ 2.2 Solvers of dynamic system
The governing equations of nonlinear dynamic systems are rarely solvable in closed form, and thus numerical integration schemes are commonly employed to approximate their time evolution. Classical approaches include implicit and explicit methods such as the Newmark-$\beta$ method, widely used in structural dynamics [newmark1959method], linear multistep methods including Adams-Bashforth and Adams-Moulton schemes [butcher2016numerical], and the family of Runge-Kutta algorithms [iserles2009first]. Each solver provides a trade-off between accuracy, stability, and computational efficiency, and the choice of method is usually problem dependent.
For hysteretic models such as the Bouc-Wen formulation, the dynamics are described by a set of coupled nonlinear ordinary differential equations (ODEs) involving both the displacement response and an internal hysteretic variable. These equations can be integrated using standard time-stepping schemes, provided that sufficient resolution is maintained to capture the nonlinear hysteretic transition. In practice, explicit solvers are often favored for their simplicity and efficiency when the system stiffness is moderate.
Among these available options, the classical fourth-order Runge-Kutta (RK4) method is particularly attractive, offering a good balance between accuracy and computational cost. It is easy to implement and has been widely applied in physics and engineering for nonlinear system simulations. In this research, the RK4 scheme is embedded into the internal variable learning process to integrate the hysteretic dynamics, yielding reliable reference trajectories that facilitate subsequent equation discovery.
@@ 2.3 Symbolic regression
Symbolic regression (SR) is a data-driven modeling approach that seeks to uncover the underlying mathematical relationships governing a system by simultaneously identifying both the model structure and its parameters [cranmer2023interpretable]. Unlike conventional regression methods that assume a predefined functional form and focus solely on parameter estimation, SR performs a global search over the space of possible mathematical expressions – typically represented as expression trees – to find the model that best fits the data. This search is guided by principles of accuracy, parsimony, and interpretability.
@@
$y_{i}\approx h(\mathbf{x}_{i};\boldsymbol{\theta}),\quad\forall i=1,\dots,N,$ (5)
where $\mathcal{H}$ denotes the space of candidate symbolic expressions constructed from a predefined set of mathematical operators (e.g., $\times,\div,\sin,\cos$...), and $\boldsymbol{\theta}$ represents any tunable parameters involved in the expression. The optimal expression $h^{*}$ is typically obtained by solving:
@@
A defining feature of SR is its remarkable ability to discover explicit closed-form expressions without requiring prior specification of the model structure [quade2016prediction]. This property is particularly desirable in scientific discovery tasks where transparency and analytical interpretability are truly essential. Unlike black-box models such as neural networks, SR yields interpretable equations that can be directly analyzed, validated, or modified by domain experts. Recent algorithmic advances, including genetic programming, Monte Carlo tree search, and hybrid gradient-evolution strategies, have substantially improved both the computational efficiency and the overall expression quality of SR-generated models. Among these, notably, PySR [cranmer2023interpretable] integrates evolutionary search with just-in-time compiled operators, thereby enabling scalable and efficient exploration of large expression spaces in scientific applications. Its application within the proposed framework is described in Section 3.3.
@@ 3 Methodology
Figure 2: The proposed unified framework for equation discovery and dynamic prediction of hysteretic systems. Problem definition and formulation: The state-space form and learnable hysteretic model are established to describe the system dynamics. Two cases in Figure 1 are considered as typical examples, where the hysteretic link equation ($g_{\phi}$) is unknown. The internal hysteretic variable $z$ is treated as a learnable parameter and jointly optimized during training. Its time derivative is computed via finite differences, denoted Diff(z)(z). Step 1: Small-amplitude responses are employed to estimate the initial parameter values. A differentiable ODE solver is embedded to learn internal variable $z$ and system parameters, with the selection matrix $\mathbf{S}$ (as in Fig. 1) specifying the observed components in the loss function. Step 2: Symbolic regression (SR) extracts explicit and interpretable governing equations for both the dynamics motion equation ($f_{\theta}$) (if Case 2) and hysteretic link equation ($g_{\phi}$).
@@ 3.1 Problem definition and formulation
In this research, our objective is to output explicit differential equation representations of hysteretic systems, with both equation discovery and dynamic prediction. The proposed framework first introduces a state–space formulation (detailed in Eq.(7)), as illustrated in Figures 1 and 2, where $f_{\theta}$ denotes the dynamic motion equation and $g_{\phi}$ represents the hysteretic link equation. The proposed unified framework is applicable to all four cases shown in Figure 1. Here, we focus on two more complex cases in which the hysteretic link equation ($g_{\phi}$) is unknown, namely, Case 1: Hysteresis Discovery and Case 2: Full Equation Discovery. The remaining two cases can be addressed using analogous procedures within the same framework. Accordingly, our goal is to identify the explicit expressions of both $f_{\theta}$ and $g_{\phi}$.
The proposed framework then introduces a learnable hysteretic model and provides the corresponding formulations for different cases. The key distinction between Case 1 and Case 2 lies in whether the structure of the dynamic motion equation ($f_{\theta}$) is known. For Case 1: Hysteresis Discovery, assuming that the dynamic motion equation is known as in Eq.(3), we can get the expression of the acceleration term $\ddot{x}$. That is, $f_{\theta}$ can be expressed with unknown parameters $m,c,k,\alpha$ as $f_{\theta}=(u-c\dot{x}-kx-\alpha z)/m$ ($t$ is omitted for brevity). Here, it is also equivalent to $f_{\theta}$ where $\theta=\{m,c,k,\alpha\}$, and $z$ is treated as a learnable parameter. For Case 2: Full Equation Discovery, since the structure of the dynamic motion equation is unknown, it is represented by a neural network $\mathcal{N}(\cdot)$, parameterized by learnable parameters $\theta$, i.e., $f_{\theta}=\mathcal{N}(x,\dot{x},u;\theta,z)$. Here, the internal variable $z$ is also treated as an additional learnable parameter within the model.
In both Case 1 and Case 2, the structure of the hysteretic link equation ($g_{\phi}$) remains unknown. Therefore, the internal hysteretic variable $z$ is regarded as a learnable parameter and jointly optimized during training. Its time derivative is computed using finite differences, denoted as Diff(z)(z). Unlike the neural network representation adopted for $f_{\theta}$ in Case 2: Full Equation Discovery, $z$ is directly modeled as a learnable parameter rather than through another neural network. This design choice is adopted since the hysteretic link ($g_{\phi}$) has a more complex expression that is difficult to represent accurately using a neural network, while only the value of $z$ is required in this step. Therefore, treating $z$ as a learnable parameter provides a more flexible and accurate formulation. Accordingly, the hysteretic link equation can be expressed as $g_{\phi}=\text{Diff}(z)$, which maintains consistency with the overall problem formulation. Thus, the formulations corresponding to both cases are established. For the other two cases illustrated in Figure 1, where the hysteretic link equation is known (e.g., Eq.(4)), the same formulation strategy can be readily applied. With these formulations in place, the subsequent steps—Solver-based internal variable learning and Equation discovery—can then be performed conveniently.
3.2 Solver-based internal variable learning

### v2

@@ Abstract
Hysteresis is a nonlinear phenomenon with memory effects, where a system’s output depends on both its current state and past states. It is prevalent in various physical and mechanical systems, such as yielding structures under seismic excitation, ferromagnetic materials, and piezoelectric actuators. Analytical models like the Bouc-Wen model are often employed but rely on idealized assumptions and careful parameter calibration, limiting their applicability to diverse or mechanism-unknown behaviors. Existing equation discovery approaches for hysteresis are often system-specific or rely on predefined model libraries, which limit their flexibility and ability to capture the hidden mechanisms.
@@
Hysteretic dynamic systems; equation discovery; symbolic regression; hysteretic variable learning
@@ 1 Introduction
Hysteresis is a common nonlinear phenomenon observed in a broad spectrum of engineering and physical systems, including ferromagnetic materials [20], piezoelectric actuators [1], structural materials [3], damping devices [22, 15], and shape memory alloys [33]. Hysteresis, characterized by its path-dependent behavior, refers to systems in which the output depends not only on the current state but also its history, often appearing as input-output loops associated with energy dissipation and irreversible processes [43, 40]. In nonlinear dynamic systems, hysteresis presents significant challenges for modeling, control, and prediction, owing to its multivalued behavior, rate dependence, and sensitivity to input history. Accurately capturing hysteresis behavior is therefore crucial in applications such as structural health monitoring [11], smart material design [9], and precision actuation [29, 41].
In response to these needs, numerous mathematical models have been proposed over the past decades to characterize hysteretic behaviors. Among these, the Bouc-Wen model [43] has been widely used as a baseline in structural dynamics due to its compact differential representation and the ability of its parameter set to reproduce a range of smooth, rate-dependent hysteresis loops. Its parameters provide physical interpretability and offer flexibility in capturing both softening and hardening behaviors.
@@
In recent years, data-driven modeling has opened new venues for discovering governing laws directly from measured data [38, 12, 16, 47, 46, 45, 48], The sparse identification of nonlinear dynamics (SINDy) [6] has shown strong potential in extracting compact and interpretable dynamic equations, by applying sparse regression with a predefined library of candidate functions. Extensions such as SINDy with control [6], physics-informed SINDy [8, 13, 25], and a sparse structural system identification method applied to hysteretic systems [23] have significantly expanded its applicability to control systems, hidden-variable models, constrained physical systems, and hysteretic dynamics (our focus in this paper). Nevertheless, the effectiveness of SINDy-type methods strongly relies on the expressiveness of the chosen function library, which constrains their ability to capture non-polynomial dynamics, rate-dependent behaviors, and discontinuities. Selecting an appropriate function library is itself a nontrivial task, often requiring prior knowledge or trial-and-error tuning.
@@
Figure 1: Classification of equation discovery approaches for hysteretic systems according to the assumed knowledge of the primary dynamics motion equation ($f_{\theta}$) and the hysteretic link equation ($g_{\phi}$). All these variables are dependent on $t$, which is omitted here for brevity. Each quadrant represents a level of prior structural specification.
@@
A variety of approaches have been proposed for equation discovery in hysteretic systems, which in this paper are classified according to the extent of prior knowledge assumed about the governing equations. In Figure 1, we organize the literature into four quadrants based on whether the structure of the primary dynamics equation ($f_{\theta}$) and the hysteretic link equation ($g_{\phi}$) are explicitly known or unknown. In the majority existing studies [23, 4], structural dynamics is assumed to follow an explicit known forms, while the hysteretic component is represented by an additional differential equation whose functional forms are to be discovered. Under this setting, the equation discovery task is reduced to symbolically identifying the hysteretic link equation (Hysteresis Discovery quadrant) [42, 28] or estimating parameters when structures of both equations are known (Parameter Estimation quadrant) [24, 44, 10]. Such assumptions greatly simplify the problem but restrict applicability to systems where the underlying structural model is already well characterized. On the other hand, frameworks that assume unknown structures in the primary dynamics equation (Dynamics Discovery quadrant) are far less common, particularly when hysteretic effects are present. When both the structural dynamics and the hysteretic link equations are unknown (Full Equation Discovery quadrant), the problem becomes substantially more challenging, requiring the simultaneous inference of two coupled, potentially nonlinear and memory-dependent governing equations.
@@
In this research, we propose a unified framework that integrates hysteretic variable learning and symbolic regression (SR) to address the challenges of hysteretic equation discovery. The framework enables direct inference of hysteretic variables and recovery of explicit and interpretable governing equations without relying on predefined model libraries. Our contributions are as follows: (i) We provide a classification of equation discovery approaches for hysteretic systems and establish a unified framework that accommodates different levels of prior knowledge; (iii) We validate the framework on synthetic benchmarks, a complex structure with high-order nonlinear and fractional terms, shake-table experiments, , demonstrating superior accuracy and generalizability compared with existing methods.
@@ 2.1 Hysteretic models
In modeling hysteretic systems, hysteretic variables are commonly introduced to mathematically represent memory effects, history-dependent behavior, which allows the system’s response to account for past states.
Among hysteresis models, the Bouc-Wen model [43] is one of the most widely used, owing to its simplicity, and its capacity to represent a broad range of hysteretic responses. It provides a set of differential equations to describe the evolution of hysteretic variables, and is capable of capturing both rate-dependent and asymmetric behaviors. The dynamics motion equation for a single degree-of-freedom (SDOF) nonlinear mass-spring-damper system with hysteresis can be expressed as follows:
@@
where $k$ is the stiffness, $\alpha$ is the hysteretic coefficient, and $z(t)$ is the hysteretic variable. Substituting Eq.(2) into Eq.(1), the equation of motion is rewritten as:
@@
The evolution of the hysteretic variable $z(t)$ is governed by the Bouc-Wen hysteresis differential equation, which can also be written as an explicit function of $\dot{x}(t)$ and $z(t)$ for compactness and interpretability (for conciseness, we drop $t$ in the equation):
@@
In this formulation, the parameter $A$ serves as a linear amplification factor that scales the influence of the displacement rate on the evolution of the hysteretic variable. The coefficients $\beta$ and $\gamma$ jointly shape the energy dissipation and curvature of the hysteresis loop, with $\beta$ primarily influencing the width and dissipation of the loop, while $\gamma$ controls its nonlinear curvature. In most existing research [36, 23, 28], the exponent $n$ is typically restricted to integer values. In this research, we generalize the formulation by allowing $n$ to take more general, including fractional values, thereby enabling richer characterization of various hysteresis.
@@ 2.2 Solvers of dynamic system
The governing equations of nonlinear dynamic systems are rarely solvable in closed form, and thus numerical integration schemes are commonly employed to approximate their time evolution. Classical approaches include implicit and explicit methods such as the Newmark-$\beta$ method, widely used in structural dynamics [31], linear multistep methods including Adams-Bashforth and Adams-Moulton schemes [7], and the family of Runge-Kutta algorithms [18]. Each solver provides a trade-off between accuracy, stability, and computational efficiency, and the choice of method is usually problem dependent.
For hysteretic models such as the Bouc-Wen formulation, the dynamics are described by a set of coupled nonlinear ordinary differential equations (ODEs) involving both the displacement response and an hysteretic variable. These equations can be integrated using standard time-stepping schemes, provided that sufficient resolution is maintained to capture the nonlinear hysteretic transition. In practice, explicit solvers are often favored for their simplicity and efficiency when the system stiffness is moderate.
Among these available options, the classical fourth-order Runge-Kutta (RK4) method is particularly attractive, offering a good balance between accuracy and computational cost. It is easy to implement and has been widely applied in physics and engineering for nonlinear system simulations. In this research, the RK4 scheme is embedded into the hysteretic variable learning process to integrate the hysteretic dynamics, yielding reliable reference trajectories that facilitate subsequent equation discovery.
@@ 2.3 Symbolic regression
Symbolic regression (SR) is a data-driven modeling approach that seeks to uncover the underlying mathematical relationships governing a system by simultaneously identifying both the model structure and its parameters [14]. Unlike conventional regression methods, SR does not assume a predefined functional form and focuses solely on parameter estimation. SR performs a global search over the space of possible mathematical expressions – typically represented as expression trees – to find the model that best fits the data. This search is guided by principles of accuracy, parsimony, and interpretability.
@@
$y_{i}\approx h(\mathbf{x}_{i};\delta),\quad\forall i=1,\dots,N,$ (5)
where $\mathcal{H}$ denotes the space of candidate symbolic expressions constructed from a predefined set of mathematical operators (e.g., $\times,\div,\sin,\cos$...), and $\delta$ represents any tunable parameters involved in the expression. The optimal expression $h^{*}$ is typically obtained by solving:
@@
A feature of SR is its ability to discover explicit closed-form expressions without requiring prior specification of the model structure [37]. This property is particularly desirable in scientific discovery tasks where transparency and analytical interpretability are truly essential. Unlike black-box models such as neural networks, SR yields interpretable equations that can be directly analyzed, validated, or modified by domain experts. Recent algorithmic advances, including genetic programming, Monte Carlo tree search, and hybrid gradient-evolution strategies, have substantially improved both the computational efficiency and the overall expression quality of SR-generated models. Among these, notably, PySR [14] integrates evolutionary search with just-in-time compiled operators, thereby enabling scalable and efficient exploration of large expression spaces in scientific applications.
@@ 3 Methodology
Figure 2: The proposed unified framework for equation discovery and dynamic prediction of hysteretic systems. Problem definition and formulation: Two cases in Figure 1 are considered as typical examples, where the hysteretic link equation ($g_{\phi}$) is unknown. Step 1: Small-amplitude responses are employed to . with the selection matrix $\mathbf{S}$ (as in Figure 1) specifying the observed components in the loss function. Step 2: Symbolic regression (SR) extracts explicit and interpretable governing equations for both the dynamics motion equation ($f_{\theta}$) (if Case 2) and hysteretic link equation ($g_{\phi}$).
@@ 3.1 Problem definition and formulation
In this research, our objective is to output explicit differential equation representations of hysteretic systems, with both equation discovery and dynamic prediction. The proposed framework first introduces a state–space formulation:
$\frac{d}{dt}\begin{bmatrix}x\\ \dot{x}\\ z\end{bmatrix}=\begin{bmatrix}\dot{x}\\ f_{\theta}(x,\,\dot{x},\,u,\,z)\\ g_{\phi}\end{bmatrix}$, (7)
@@
where $x$ denotes the displacement, and $\dot{x}$ is the velocity. $f_{\theta}$ denotes the dynamic motion equation and $g_{\phi}$ represents the hysteretic link equation. The proposed unified framework is applicable to all four cases shown in Figure 1. Here, we focus on two more complex cases in which $g_{\phi}$ is unknown, namely, Case 1: Hysteresis Discovery and Case 2: Full Equation Discovery. The remaining two cases can be addressed using analogous procedures within the same framework.
3.2 Solver-based hysteretic variable learning